# Prediction of two-neutron halos in the $N=28$ isotones ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$

### v1

@@ IV Three-body ($\text{core}+n+n$) calculations
Using the core$+n$ potentials outlined in the preceding section, we compute the three-body ground-state of the ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ nuclei assuming a simple description in which two valence neutrons couple to $J^{\pi}=0^{+}$. To provide a more comprehensive description, it is essential to have precise experimental data on the core$+n$ spectrum and a thorough understanding of the intricate core$+n$ interactions that can result in the splitting of single-particle energy levels. In the present calculation, we also need a prescription for the $nn$ interaction. We use the Gogny-Pires-Tourreil (GPT) potential Gogny et al. (1970) as in previous works Lovell et al. (2017); Thompson et al. (2004); Casal et al. (2013). It is important to emphasize that when it comes to the ground-state properties of $\text{core}+n+n$ systems, the specific selection of the $nn$ interaction is not particularly critical, as long as the chosen interaction adequately represents $nn$ scattering data, as discussed in Ref. Zhukov et al. (1993). As previously introduced in Eq. (5), we incorporate a phenomenological three-body force to obtain the energy of the three-body ground state, following a methodology similar to that employed in our earlier studies Casal et al. (2013); Singh et al. (2020, 2022). The three-body problem is solved by diagonalizing the Hamiltonian in a THO basis including hyperspherical harmonics up to a maximum hypermomentum $K_{max}$ in the wave-function expansion (Eq. (1)), and $i=0,\dots,N$ excitations for the hyperradial functions (Eq. (3)). Note that fixing $K_{max}$ implies that the orbital angular momenta associated with each Jacobi coordinate are restricted to $l_{x}+l_{y}\leq K$ Zhukov et al. (1993). In this work, $K_{max}=46$ and $N=30$ are found to provide converged results.
@@
In order to compute the ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ matter radius within the present three-body model Casal (2016) using $R_{m}=\sqrt{\frac{1}{A}\left(A_{c}R_{c}^{2}+\langle r_{nn}^{2}\rangle\right)}$, the size of the ${}^{38}\mathrm{Mg}$ and ${}^{37}\mathrm{Na}$ cores are required as input. We have used $R_{c}=R_{m}(^{38}\mathrm{Mg},^{37}\mathrm{Na})=3.600,3.637$ fm, respectively, adopted from Ref. Watanabe et al. (2014); Geng et al. (2004). These are somewhat uncertain, so it is more precise to look at a comparison of the differences across various potential models as well as the relative increase between core and three-body systems.
@@
[…]
The computed partial wave content and radii for each $s_{2n}$ value are shown in the upper panel and lower panel, respectively, of Fig 2. As expected, shallower (bound) ground states yield larger $p$-wave occupancy which in turn leads to an “inverted” picture, in contrast to the deeply-bound ground states which lead to the “degenerate” and “normal” pictures for all sets. As can be seen from the upper panel of Fig. 2, when moving from Set 1 to Set 3 the $p$-wave content in the ground state increases, whereas the $f$-wave content decreases for each choice of $s_{2n}$ value. We remark that a similar behavior was found in Ref.Masui et al. (2020).
@@
In the lower panel of Fig 2, we show the radius of the ${}^{38}\mathrm{Mg}$ and ${}^{37}\mathrm{Na}$ cores as a horizontal dotted black line which makes it easy to extract the relative change in matter radius of ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ relative to the radius of the appropriate core, $\Delta R=R_{m}-R_{core}$. With the current uncertainties in $s_{2n}$ and the low-lying spectrum of subsystems, the difference of the matter radius of these nuclei and their cores ranges between $0.1$-$0.5$ fm for different choices of potential sets and $s_{2n}$ values. This number is notably smaller than those for well-established halo nuclei such as ${}^{6}\mathrm{He}$ or ${}^{11}\mathrm{Li}$, but is close to the heaviest known two-neutron halo, ${}^{29}\mathrm{F}$, which is only $0.35$ fm larger than the ${}^{27}\mathrm{F}$ core Singh et al. (2020); Casal et al. (2020); Bagchi et al. (2020). It is worth mentioning that, recently, we have shown that the halo formation in ${}^{29}\mathrm{F}$ is connected to the weakening of the $N=20$ shell gap, which leads to the intrusion of the $p_{3/2}$ orbital and reduced binding Singh et al. (2020); Fortunato et al. (2020); Casal et al. (2020), which leads to halo formation. The same scenario is observed when $N=28$ melts, with the difference that the $p_{3/2}$ orbital is within the same major shell. As depicted in Fig.2, our three-body findings reveal a substantial increase in $\Delta R$ for the least deeply bound state, with a substantial $p_{3/2}$ component. This suggests that our study provides evidence for a halo structure in the ground states of both ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$. However, experimental cross section measurements are necessary to confirm this assertion.
Figure 3: The Variation of matter radii ($R_{m}$) for Mg (upper panel) and Na (lower panel) isotopes with mass number $A$. Experimental values, shown in black, blue, and red circles, were taken from Refs. Ozawa et al. (2001), Kanungo et al. (2011), and Takechi et al. (2014), respectively. The dotted black line corresponds to the $R_{0}A^{1/3}$ fit for $A=24$-$37$ for Mg-isotopes and $A=23$-$32$ for Na-isotopes. It should be noted that each colour band associated with a specific $s_{2n}$ corresponds to the range of values for Sets $1$, $2$, and $3$, and the same applies to Fig.  5.
@@
V Total reaction cross section
@@
Experimentally, a very obvious way to determine whether a nucleus is a halo nucleus, is to look for an enhanced reaction cross section. Thus we examine the total reaction cross section ($\sigma_{R}$) by employing the conventional Glauber theory Glauber (1959). We utilize the nucleon-target formalism Abu-Ibrahim and Suzuki (2000) and apply the nucleon-nucleon profile function from Refs. Abu-Ibrahim et al. (2008, 2009, 2010).This approach has proven effective in various high-energy nucleus-nucleus collision reactions, particularly those with unstable nuclei, and it successfully replicated isotope dependence of the total reaction cross sections with appropriate density distributions Horiuchi and Suzuki (2006); Horiuchi et al. (2007a, b, 2010b, 2012, 2015); Nagahisa and Horiuchi (2018).
@@
Figure 5: Measured and calculated total reaction cross section for different isotopes of Mg and Na at different energies. Theoretical results (shown with green, magenta, and indigo bars) for ${}^{40}\mathrm{Mg}$$+$${}^{12}\mathrm{C}$ are shown at (a) $240$ MeV/A and (b) $1000$ MeV/A. Experimental data shown with red, black, and blue dots are taken from Takechi et al. (2014) (at $240$ MeV/A), Kanungo et al. (2011) (at $900$ MeV/A), and Ozawa et al. (2001) (at $950$ MeV/A), respectively. (c) Shows the results for ${}^{39}\mathrm{Na}$$+$${}^{12}\mathrm{C}$ at $1000$ MeV/A and experimental data shown with black dots is taken from Ozawa et al. (2001) (at $950$ MeV/A).
@@
The additional theoretical inputs required for this reaction model include the density distributions of both the projectile and target nuclei. In the case of ${}^{40}\mathrm{Mg}$, we first generate a harmonic-oscillator (HO) type density distribution for ${}^{38}\mathrm{Mg}$ that reproduces the measured total reaction cross section of ${}^{38}\mathrm{Mg}$$+$${}^{12}\mathrm{C}$ by Takechi et al., Takechi et al. (2014) at $240$MeV/nucleon ($1535\pm 21$ mb) by assuming a simple shell model state $(1f_{7/2})_{\nu}^{8}\otimes(1d_{5/2})_{\pi}^{4}$. The density for ${}^{40}\mathrm{Mg}$ was constructed by simply adding the $2n$ densities calculated in the previous section to the ${}^{38}\mathrm{Mg}$ density. No center-of-mass correction was applied in these calculations, but we believe this is a safe strategy for a system with such a heavy core. Using this prescription, we predict the $\sigma_{R}$ for ${}^{40}\mathrm{Mg}$ at different incident energies, $240$ MeV/nucleon (lies between $1601$-$1807$ mb, shown in Fig. 5(a)) and $1000$ MeV/A (lies between $1695$-$1944$ mb, shown in the Fig. 5(b)). As can be seen from Fig. 5(a) and (b), the predicted values of $\sigma_{R}$ for ${}^{40}\mathrm{Mg}$ show significant enhancement with respect to the observed $\sigma_{R}$ in the lower-$A$ isotopes for both choices of energy. This shows that we can use these reaction cross sections as an indication of the melting of the $N=28$ shell closure.
@@
We use the same HO parameterization established for ${}^{38}\mathrm{Mg}$ for ${}^{37}\mathrm{Na}$. Hence, we forecast the $\sigma_{R}$ for ${}^{39}\mathrm{Na}$ at an energy of $1000$ MeV/nucleon, as depicted in Fig. 5(c) (lies between $1710$-$1894$ mb). Our computed estimates of $\sigma_{R}$ for ${}^{39}\mathrm{Mg}$ reveal a noteworthy increase compared to the observed $\sigma_{R}$ in its lower-mass isotopes. This reinforces our earlier finding in the ${}^{40}\mathrm{Mg}$, which points to the formation of a two-neutron halo in the ground state of ${}^{39}\mathrm{Na}$, leading to the disruption of the neutron shell at $N=28$.
VI Conclusions
@@
The matter radii of the ${}^{38}\mathrm{Mg}$ and ${}^{37}\mathrm{Na}$ cores exhibit a relative increase falling within the range of $0.1$ to $0.5$ fm in various scenarios, lending support to the idea of a halo structure in the ground state of ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$. This assertion gains further backing from the elevated predicted reaction cross sections for these nuclei in comparison to the lower mass isotopes. Nevertheless, it is imperative to verify this conclusion through experimental measurements of interaction cross sections. Future perspectives include incorporating core deformation within the effective three-body description explicitly, to assess its influence on the halo structure in this region of the nuclear chart.
@@ Acknowledgments
This work was supported by the UK Science and Technology Funding Council [grant number ST/V001116/1] (JS and NRW), by the Polish National Science Centre (NCN) [under Contract No 2018/31/B/ST2/02220] (JS and WS), by the Spanish MCIN/AEI/10.13039/501100011033 [under I+D+i project No. PID2020-114687GB-I00] (JC), and by the JSPS KAKENHI Grants Nos. 18K03635 and 22H01214 (WH).

### v2

@@ 4 Three-body ($\text{core}+n+n$) calculations
Using the core$+n$ potentials outlined in the preceding section, we compute the three-body ground-state of the ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ nuclei assuming a simple description in which two valence neutrons couple to $J^{\pi}=0^{+}$. To provide a more comprehensive description, it is essential to have precise experimental data on the core$+n$ spectrum and a thorough understanding of the intricate core$+n$ interactions that can result in the splitting of single-particle energy levels. In the present calculation, we also need a prescription for the $nn$ interaction. We use the Gogny-Pires-Tourreil (GPT) potential [67] as in previous works [68, 56, 54]. It is important to emphasize that when it comes to the ground-state properties of $\text{core}+n+n$ systems, the specific selection of the $nn$ interaction is not particularly critical, as long as the chosen interaction adequately represents $nn$ scattering data, as discussed in Ref. [48]. As previously introduced in Eq. (5), we incorporate a phenomenological three-body force to obtain the energy of the three-body ground state, following a methodology similar to that employed in our earlier studies [54, 30, 58]. The three-body problem is solved by diagonalizing the Hamiltonian in a THO basis including hyperspherical harmonics up to a maximum hypermomentum $K_{max}$ in the wave-function expansion (Eq. (1)), and $i=0,\dots,N$ excitations for the hyperradial functions (Eq. (3)). Note that fixing $K_{max}$ implies that the orbital angular momenta associated with each Jacobi coordinate are restricted to $l_{x}+l_{y}\leq K$ [48]. In this work, $K_{max}=46$ and $N=30$ are found to provide converged results.
@@
In order to compute the ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ matter radius within the present three-body model [55] using $R_{m}=\sqrt{\frac{1}{A}\left(A_{c}R_{c}^{2}+\langle r_{nn}^{2}\rangle\right)}$, the size of the ${}^{38}\mathrm{Mg}$ and ${}^{37}\mathrm{Na}$ cores are required as input. We have used $R_{c}=R_{m}(^{38}\mathrm{Mg},^{37}\mathrm{Na})=3.600,3.637$ fm, respectively, adopted from Ref. [69, 70]. These are somewhat uncertain, so it is more precise to look at a comparison of the differences across various potential models as well as the relative increase between core and three-body systems.
Figure 2: (upper panel) The occupancy of the $(f_{7/2})^{2}$ (solid lines) and $(p_{3/2})^{2}$ orbits (dashed lines) for different parameter sets in ${}^{40}\mathrm{Mg}$ (left) and ${}^{39}\mathrm{Na}$ (right), respectively, as a function of $s_{2n}$. (lower panel) Variation of matter radii (Rm) for ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ with $s_{2n}$. The black dotted line shows the core radius, ${}^{38}\mathrm{Mg}$ and ${}^{37}\mathrm{Na}$, respectively. It should be noted that the labels Inverted, Degenerate and Normal pictures here refer to the outcome of full three-body calculations and should not be confused with the single-particle scenarios of Table 1.
@@
The computed partial wave content and radii for each $s_{2n}$ value are shown in the upper panel and lower panel, respectively, of Fig 2. As expected, shallower (bound) ground states yield larger $p$-wave occupancy which in turn leads to an “inverted” picture, in contrast to the deeply-bound ground states which lead to the “degenerate” and “normal” pictures for all sets. As can be seen from the upper panel of Fig. 2, when moving from Set 1 to Set 3 the $p$-wave content in the ground state increases, whereas the $f$-wave content decreases for each choice of $s_{2n}$ value. We remark that the similar behavior was found in Ref.[71].
@@
In the lower panel of Fig 2, we show the radius of the ${}^{38}\mathrm{Mg}$ and ${}^{37}\mathrm{Na}$ cores as a horizontal dotted black line which makes it easy to extract the relative change in matter radius of ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ relative to the radius of the appropriate core, $\Delta R=R_{m}-R_{core}$. With the current uncertainties in $s_{2n}$ and the low-lying spectrum of subsystems, the difference of the matter radius of these nuclei and their cores ranges between $0.1$-$0.5$ fm for different choices of potential sets and $s_{2n}$ values. This number is notably smaller than those for well-established halo nuclei such as ${}^{6}\mathrm{He}$ or ${}^{11}\mathrm{Li}$, but is close to the heaviest known two-neutron halo, ${}^{29}\mathrm{F}$, which is only $0.35$ fm larger than the ${}^{27}\mathrm{F}$ core [30, 32, 13]. It is worth mentioning that, recently, we have shown that the halo formation in ${}^{29}\mathrm{F}$ is connected to the weakening of the $N=20$ shell gap, which leads to the intrusion of the $p_{3/2}$ orbital and reduced binding [30, 31, 32], which leads to halo formation. The same situation is observed when $N=28$ melts, with the difference that the $p_{3/2}$ orbital is within the same major shell. As depicted in Fig.2, our three-body findings reveal a substantial increase in $\Delta R$ for the least deeply bound state, with a substantial $p_{3/2}$ component. This suggests that our study provides evidence for a halo structure in the ground states of both ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$. However, experimental cross section measurements are necessary to confirm this assertion.
Figure 3: The Variation of matter radii ($R_{m}$) for Mg (upper panel) and Na (lower panel) isotopes with mass number $A$. Experimental values, shown in black, blue, and red circles, were taken from Refs. [72], [73], and [34], respectively. The dotted black line corresponds to the $R_{0}A^{1/3}$ fit for $A=24$-$37$ for Mg-isotopes and $A=23$-$32$ for Na-isotopes. It should be noted that each colour band associated with a specific $s_{2n}$ corresponds to the range of values for Sets $1$, $2$, and $3$, and the same applies to Fig.  5.
@@
5 Total reaction cross section
@@
Experimentally, a very obvious way to determine whether a nucleus is a halo nucleus, is to look for an enhanced reaction cross section. Thus we examine the total reaction cross section ($\sigma_{R}$) by employing the conventional Glauber theory [74]. We utilize the nucleon-target formalism [75] and apply the nucleon-nucleon profile function from Refs. [76, 77, 78].This approach has proven effective in various high-energy nucleus-nucleus collision reactions, particularly those with unstable nuclei, and it successfully replicated isotope dependence of the total reaction cross sections with appropriate density distributions [60, 79, 80, 81, 82, 83, 84].
@@
Figure 5: Measured and calculated total reaction cross section for different isotopes of Mg and Na at different energies. Theoretical results (shown with green, magenta, and indigo bars) for ${}^{40}\mathrm{Mg}$$+$${}^{12}\mathrm{C}$ are shown at (a) $240$ MeV/A and (b) $1000$ MeV/A. Experimental data shown with red, blue, and black dots are taken from [34] (at $240$ MeV/A), [73] (at $900$ MeV/A), and [72] (at $950$ MeV/A), respectively. (c) Shows the results for ${}^{39}\mathrm{Na}$$+$${}^{12}\mathrm{C}$ at $1000$ MeV/A and experimental data shown with black dots is taken from [72] (at $950$ MeV/A). The dotted black line corresponds to a $\sigma_{0}A^{2/3}$ fit to guide the eye of the reader.
@@
The additional theoretical inputs required for this reaction model include the density distributions of both the projectile and target nuclei. In the case of ${}^{40}\mathrm{Mg}$, we first generate a harmonic-oscillator (HO) type density distribution for ${}^{38}\mathrm{Mg}$ that reproduces the measured total reaction cross section of ${}^{38}\mathrm{Mg}$$+$${}^{12}\mathrm{C}$ by Takechi et al., [34] at $240$MeV/nucleon ($1535\pm 21$ mb) by assuming a simple shell model state $(1f_{7/2})_{\nu}^{8}\otimes(1d_{5/2})_{\pi}^{4}$. The density for ${}^{40}\mathrm{Mg}$ was constructed by simply adding the $2n$ densities calculated in the previous section to the ${}^{38}\mathrm{Mg}$ density. No center-of-mass correction was applied in these calculations, but we believe this is a safe strategy for a system with such a heavy core. Using this prescription, we predict the $\sigma_{R}$ for ${}^{40}\mathrm{Mg}$ at different incident energies, $240$ MeV/nucleon (lies between $1601$-$1807$ mb, shown in Fig. 5(a)) and $1000$ MeV/A (lies between $1695$-$1944$ mb, shown in the Fig. 5(b)). As can be seen from Fig. 5(a) and (b), the predicted values of $\sigma_{R}$ for ${}^{40}\mathrm{Mg}$ show significant enhancement with respect to the observed $\sigma_{R}$ in the lower-$A$ isotopes for both choices of energy. This shows that we can use these reaction cross sections as an indication of the melting of the $N=28$ shell closure.
@@
We use the same HO parameterization established for ${}^{38}\mathrm{Mg}$ for ${}^{37}\mathrm{Na}$. Hence, we forecast the $\sigma_{R}$ for ${}^{39}\mathrm{Na}$ at an energy of $1000$ MeV/nucleon, as depicted in Fig. 5(c) (lies between $1710$-$1894$ mb). Our computed estimates of $\sigma_{R}$ for ${}^{39}\mathrm{Na}$ reveal a noteworthy increase compared to the observed $\sigma_{R}$ in its lower-mass isotopes. This reinforces our earlier finding in the ${}^{40}\mathrm{Mg}$, which points to the formation of a two-neutron halo in the ground state of ${}^{39}\mathrm{Na}$, leading to the disruption of the neutron shell at $N=28$.
6 Conclusions
@@
The matter radii of ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ show an increase of $0.1$ to $0.5$ fm compared to their respective cores ${}^{38}\mathrm{Mg}$ and ${}^{37}\mathrm{Na}$, lending support to the idea of a potential halo structure in their ground state. This assertion gains further backing from the elevated predicted reaction cross sections for these nuclei in comparison to the lower mass isotopes. Nevertheless, it is imperative to verify this conclusion through experimental measurements of interaction cross sections. Future perspectives include incorporating core deformation within the effective three-body description explicitly, to assess its influence on the halo structure in this region of the nuclear chart.
@@ Acknowledgments
We thank L. Fortunato for careful reading of the manuscript. This work was supported by the UK Science and Technology Funding Council [grant number ST/V001116/1] (JS and NRW), by the Polish National Science Centre (NCN) [under Contract No 2018/31/B/ST2/02220] (JS and WS), by the Spanish MCIN/AEI/10.13039/501100011033 [under I+D+i project No. PID2020-114687GB-I00] (JC), and by the JSPS KAKENHI Grants Nos. 18K03635 and 22H01214 (WH).